A rectangular window has a perimeter of 258 inches. The width of the window is 24 inches. What is the area of the window?
Answer by TimothyLamb(4379)

p = 2L + 2w = 258
w = 24
---
2L + 2w = 258
2L + 2(24) = 258
2L + 48 = 258
2L = 210
---
L = 105
---
a = Lw
a = 105 * 24
---
a = 2520 sq.in
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php